A fabric store sells two types of ribbon. One customer buys 3 rolls of the lace ribbon and 2 rolls of the satin ribbon and has a total of 120 yards of ribbon. Another customer buys 2 rolls of the lace ribbon and 4 rolls of the satin ribbon and has a total of 160 yards of ribbon.
How many yards are on one roll of lace ribbon and one roll of satin ribbon?

Answer by ikleyn(52786)

.

    3x + 2y = 120   (1)

    2x + 4y = 160   (2)



Divide eq(2) by 2 (both sides). Keep equation (1) as is.  You will get

    3x + 2y = 120   (1')

     x + 2y =  80   (2')


Now subtract eq(2') from eq(1').  You will get

    2x      = 120 - 80 = 40

     x =               = 40/2 = 20.


Then from equation (2'),  

    20 + 2y = 80

         2y = 80 - 20 = 60

          y            = 60/2 = 30.


ANSWER.   20 yards in the lace ribbon roll and 30 yards in the satin ribbon roll.

Solved.